### 04 Mar 2026 v1

@@ 1 Introduction
Let $p(A)$ denote the expected fraction of monochromatic edges when we use a $1$-round algorithm $A$, and let $p^{*}$ be the infimum of these values over all algorithms (see Section 2 for the detailed definition). What can we say about $p^{*}$? Prior work [Hirvonen et al.(2017)] together with a trivial lower bound implies $1/5\leqslant p^{*}\leqslant 1/4$. But what is the precise value? In this work we develop new techniques for answering this question and show that
@@ 1.1 Motivation and broader context: distributed quantum advantage
While this problem is of independent interest—see e.g. [Hirvonen et al.(2017), Shearer(1992)]—our original motivation for studying it arises from the pressing open question of understanding distributed quantum advantage. One of the biggest open questions in this area boils down to whether 3-coloring cycles (or any local symmetry-breaking problem) admits a distributed quantum advantage, see e.g. [Akbari et al.(2025), Balliu et al.(2025), d’Amore(2025), Le Gall and Rosmanis(2022), Gavoille et al.(2009)]. Currently, we cannot even exclude the possibility that quantum-LOCAL can 3-color cycles in 1 round, with probability 1.
@@ 1.2 Coworkers: large language models and proof assistants
As it is hard to verify nontrivial LLM-generated mathematical proofs, we also had LLMs formalize both the upper bound and the lower bound in Lean 4 proof assistant, so that the proof can be automatically verified with a computer; see [GPT-5.2 et al.(2026)] for the full Lean 4 code.
@@ 3 Simple observations and prior work
Third, we can use the idea familiar from e.g. [Hirvonen et al.(2017)]: find a random cut and use 1 round to “improve” it. The obvious rule for improvement is this: if I am on the same side as my predecessor and successor, I will change sides. This rule $f_{3}$ results in $p(f_{3})=1/4$, which is the tightest upper bound we have seen in prior work.
@@ 4.3 Lower bound
We use a semidefinite-program (SDP) relaxation [Vandenberghe and Boyd(1996)] of the max-cut problem [Goemans and Williamson(1995)], with correlation triangle inequalities [Deza and Laurent(1997)]. Graph $\operatorname{DB}_{\operatorname{distinct}}(n)$ is highly symmetric, and by exploiting symmetries [Gatermann and Parrilo(2004)], the size of the SDP is bounded by a constant independent of $n$. We use an SDP solver to find a feasible dual solution [Peyrl and Parrilo(2008)], which can be turned into a certificate for $p_{\operatorname{distinct}}(10^{6})\geqslant 0.23879$. We refer to the formalized proof [GPT-5.2 et al.(2026)] for more details. ∎
@@ 4.4 Upper bound
We then further tuned $f^{\tau}$ and discovered that modifications shown in Fig. 3(b) lead to a family of algorithms defined with three parameters that can slightly outperform $f^{\tau}$. Optimizing these parameters results in an algorithm $f$ with $p(f)<0.24118$; again, see [GPT-5.2 et al.(2026)] for the details.

### 05 Mar 2026 v2

@@ 1 Introduction
Let $p(A)$ denote the expected fraction of monochromatic edges when we use a $1$-round algorithm $A$, and let $p^{*}$ be the infimum of these values over all algorithms (see Section 2 for the detailed definition). What can we say about $p^{*}$? Prior work [HRSS17] together with a trivial lower bound implies $1/5\leqslant p^{*}\leqslant 1/4$. But what is the precise value? In this work we develop new techniques for answering this question and show that
@@ 1.1 Motivation and broader context: distributed quantum advantage
While this problem is of independent interest—see e.g. [HRSS17, She92]—our original motivation for studying it arises from the pressing open question of understanding distributed quantum advantage. One of the biggest open questions in this area boils down to whether 3-coloring cycles (or any local symmetry-breaking problem) admits a distributed quantum advantage, see e.g. [ACRd+25, BBCR+25, d’A25, LR22, GKM09]. Currently, we cannot even exclude the possibility that quantum-LOCAL can 3-color cycles in 1 round, with probability 1.
@@ 1.2 Coworkers: large language models and proof assistants
As it is hard to verify nontrivial LLM-generated mathematical proofs, we also had LLMs formalize both the upper bound and the lower bound in Lean 4 proof assistant, so that the proof can be automatically verified with a computer; see [GS+26] for the full Lean 4 code.
@@ 3 Simple observations and prior work
Third, we can use the idea familiar from e.g. [HRSS17]: find a random cut and use 1 round to “improve” it. The obvious rule for improvement is this: if I am on the same side as my predecessor and successor, I will change sides. This rule $f_{3}$ results in $p(f_{3})=1/4$, which is the tightest upper bound we have seen in prior work.
@@ 4.3 Lower bound
We use a semidefinite-program (SDP) relaxation [VB96] of the max-cut problem [GW95], with correlation triangle inequalities [DL97]. Graph $\operatorname{DB}_{\operatorname{distinct}}(n)$ is highly symmetric, and by exploiting symmetries [GP04], the size of the SDP is bounded by a constant independent of $n$. We use an SDP solver to find a feasible dual solution [PP08], which can be turned into a certificate for $p_{\operatorname{distinct}}(10^{6})\geqslant 0.23879$. We refer to the formalized proof [GS+26] for more details. ∎
@@ 4.4 Upper bound
We then further tuned $f^{\tau}$ and discovered that modifications shown in Fig. 3(b) lead to a family of algorithms defined with three parameters that can slightly outperform $f^{\tau}$. Optimizing these parameters results in an algorithm $f$ with $p(f)<0.24118$; again, see [GS+26] for the details.